OpenStudy (guccidothedishes):

** will medal and fan What are the explicit equation and domain for an arithmetic sequence with a first term of 5 and a second term of 2?

OpenStudy (guccidothedishes):

n=5−2(n−1);allintegerswheren\[≥1 an=5−2(n−1);allintegerswheren\[≥0 an=5−3(n−1);allintegerswheren\[≥1 an=5−3(n−1);allintegerswheren\[≥0

OpenStudy (guccidothedishes):

@liljj1421 ??

OpenStudy (guccidothedishes):

@Michele_Laino

OpenStudy (michele_laino):

hint: we can write this: \[\Large \begin{gathered} {a_2} = {a_1} + d \hfill \\ 2 = 5 + d \hfill \\ \end{gathered} \] where d is the sequence constant

OpenStudy (michele_laino):

what si d?

OpenStudy (michele_laino):

after that the general formula, for the n-th term is: \[\Large {a_n} = {a_1} + \left( {n - 1} \right)d\] where a_1=5

OpenStudy (guccidothedishes):

how do I get d ?

OpenStudy (michele_laino):

it is simple, from the second equation, I can write: \[d = 2 - 5 = ...?\]

OpenStudy (guccidothedishes):

so D=3 ?

OpenStudy (michele_laino):

no, d=-3

OpenStudy (guccidothedishes):

okay so I can already eliminate 2 of the answer choices

OpenStudy (guccidothedishes):

but how do I determine the second part of the answer?

OpenStudy (michele_laino):

yes! furthermore, please keep in mind that we start to count from n=1

OpenStudy (guccidothedishes):

So my answer is C

OpenStudy (michele_laino):

yes! That's right! since if n=1, we get: \[{a_1} = {a_1} + \left( {1 - 1} \right)d = 5 + 0 \times \left( { - 3} \right) = 5\]

OpenStudy (guccidothedishes):

Thank you so much :)

OpenStudy (michele_laino):

:)